Lolwah plans to save for her 5-year doctorate degree, which starts 6 years from now. The current total expenditure is $7,200 and it is expected to grow by 7 percent when she stars her study. She identifies a savings plan that allows her to earn an interest annually of 8 percent. How much should Lolwah deposit each year, starting one year from today? (Assume that she plans to make 5 payments.)
Answer #1

Here it is assumed the rate of growth given is absolute so when she starts studying she will require 7200 * 1.06 = 7,632. She will require to deposit in 5 instalments which will amount to 7632 after growing at 8% every year.

Lolwah is investing in an annuity so she will be paying some amount every year so the formula for Future value of annuity due is

F = P * ((1+I)^N -1)/I * (1 + I)

We are taking annuity due because we are paying at the beginning of each year

Here P is the required variable while we have the Future value of Rs. 7632

7632 = P * ((1.08)^5 -1)/0.08 * 1.08

7632 = P * 5.86660096 * 1.08

P = 1204.55

If she invests 1204.55 for 5 years at 8% then at the end of 5 years she will get 7632